OpenStudy (anonymous):

perform indicated operation then put in a+bi form -i(5-i)+2i(i+6)

OpenStudy (anonymous):

-i(5-i)+2i(i+6) = -5i+i^2+2i^2+12i =3i^2 +7i =i(3i+7)